3) A tour group split into two groups when waiting in line for food at a fast
food counter. The first group bought 8 slices of pizza and 4 soft drinks
for $36.12. The second group bought 6 slices of pizza and 6 soft drinks
for $31.74. How much does one slice of pizza cost?

Answer by Cromlix(4381)

x = slices of pizza: y = soft drinks
8x + 4y = 3612....1
6x + 6y = 3174....2
multiply (1) by 6
and multiply (2) by 4
48x + 24y = 21672....3
24x + 24y = 12696....4
Subtract (4) from (3)
24x = 8976
x = 3.74
Substitute x = 374 into (1)
8x + 4y = 3612....1
8(374) + 4y = 3612
2992 + 4y = 3612
4y = 3612 - 2992
4y = 620
y = 155
A slice of pizza cost $3.74
A soft drink cost $1.55
Hope this helps
:-)